# Topological Order in the Spectral Riemann Surfaces of Non-Hermitian Systems

### v1

@@ Abstract
The topological protection of the Fermi cuts after threading and merging the EPs can be understood from the perspective of eigenenergy braids. These braids are defined on closed lines around the holes of the toroidal Brillouin zone, cf. the bottom row in Fig. 2(b). A Fermi cut results in a crossing of the braids orthogonal to the cut. This is the only possible non-trivial braid, since any pair of Fermi cuts along the same direction reduces to Fermi arcs, which have a trivial braid structure. Perturbations cannot change the non-trivial braids of single Fermi cuts, ensuring their stability. This is similar to the braid protection of third-order EPs observed in Ref. [29].
The presence or absence of non-trivially closed Fermi cuts defines four distinct states of the system. These states are only well-defined without EPs present. Tuning the system from one of these states to another thus requires the merger of all EPs after the threading procedure. To distinguish the states we define two topological invariants, which are the crossing numbers of the braids along non-contractible loops through the Brillouin zone. We associate the invariant $m_{\alpha}=0$ with the trivial braid and $m_{\alpha}=1$ with the non-trivial braid, where $\alpha\in\{x,y\}$ indicates whether the path runs along the $k_{x}$ or $k_{y}$ direction. This establishes $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ topological order in the non-Hermitian spectrum. Based on the four states distinguished by the combinations of topologically protected closed Fermi cuts, we construct the analogy to the ground-state manifold of the toric code in an exemplary model. Any closed Fermi cut is mapped to a non-contractible loop of flipped spins in the toric code, and the four different ground states are represented by $(m_{x},m_{y})\in\{(0,0),(1,0),(0,1),(1,1)\}$. The ground-state manifold is realized by the tunable non-Hermitian system with the Bloch Hamiltonian
@@
At the same time, added bands affect the ground-state manifold of the non-Hermitian multi-band model, which comprises any fully non-degenerate spectral structure. The number of topologically distinguishable Riemann surface structures increases, because additional non-contractible Fermi cuts along the $k_{x}$ and $k_{y}$ direction are possible between any two bands. However, intersecting Fermi cuts induce EP3 excitations if one energy sheet is part of both cuts. Therefore not all combinations of Fermi cuts are part of the ground-state manifold. Furthermore, Fermi cuts along the same direction and between the same energy sheets still reduce to Fermi arcs, so that one only needs to consider the presence or absence of each type of cut. A topological invariant can be defined for each ground state by assigning a color to each energy sheet, which colors the braids along non-contractible loops through the Brillouin zone. Invariants then correspond to the crossing numbers of braids in the colored (pure) braid group $PB_{n}$. This establishes the ground-state manifold as a subspace of the space with $\left(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\right)^{m}$ topological order, where $m=\sum_{j=1}^{n-1}j=\tfrac{1}{2}(n-1)n$.
@@
Implementation on metasurfaces and in single-photon interferometry.—The implementation of different topologically protected states in the non-Hermitian toric-code analog requires a high degree of parametric tunability over a toroidal parameter space. A prominent platform providing such tunability are optical or plasmonic metasurfaces, which rely on artificial units cells defined in periodic arrays of nanoantennas [34, 35]. Loss generates non-Hermitian contributions to the dynamics of these systems. By carefully adjusting the shape, spacing and orientation of the antennas in the surrounding medium or on the host surface, the parameters of the system can be precisely controlled. Such setups are capable of realizing the two-band model $H_{\text{FC}}(\bm{k})$ in a two-orbital square lattice given full control over the onsite terms and hoppings up to next-nearest-neighbor distance. Mechanical metamaterials, in which individual oscillators are driven depending on the state of the system, provide another feasible platform. Here, the dynamical control of the oscillator response allows for the observation of EP pair creation and their threading through the Brillouin zone. Thus these materials are capable of resolving the transition between topologically distinct states. Single-photon interferometry was recently established as a different platform for the study of non-Hermitian topological features [36]. These newly designed interferometers can, for example, encode the evolution of a non-Hermitian three-band Bloch Hamiltonian and realize a braid protected EP3 [36, 29]. By instead encoding the parameters of $H_{\text{FC}}$, the non-Hermitian analog of the toric-code ground states can be realized in this state-of-the-art platform.
@@
Acknowledgments.—We are grateful to Joost Slingerland for his lectures during the Young Researchers School 2024 in Maynooth. We acknowledge funding from the Max Planck Society’s Lise Meitner Excellence Program 2.0.

### v2

@@ Abstract
The topological protection of the Fermi cuts after threading and merging the EPs can be understood from the perspective of eigenenergy braids. These braids are defined on closed lines around the holes of the toroidal Brillouin zone, cf. the bottom row in Fig. 2(b). A Fermi cut results in a crossing of the braids orthogonal to the cut. Perturbations cannot change the non-trivial braids of single Fermi cuts, ensuring their stability [29]. This is similar to the braid protection of third-order EPs observed in Ref. [30].
The presence or absence of non-trivially closed Fermi cuts results in a different connection of the Riemann energy sheets. This connectedness describes the (im-)possibility to change between sheets by following a non-contractible loop $\mathcal{C}$ in the Brillouin zone. If the eigenenergy crosses an even number of Fermi cuts, it returns to its initial value after traversing the parametric loop $\mathcal{C}$–the sheets are disconnected along this direction. When crossing an odd number of Fermi cuts, the sheets are connected instead and the eigenenergy does not return to the initial value. By distinguishing whether the Riemann sheet is connected along both directions, we define four topologically distinct states of the system. These states are only well-defined without EPs present. Tuning the system from one of these states to another thus requires the merger of all EPs after the threading procedure. To distinguish the states we define two topological invariants, which are derived from the permutations of the eigenenergies along non-contractible loops through the Brillouin zone. The invariant $m_{\alpha}=0$ is associated with non-connected and $m_{\alpha}=1$ with connected Riemann sheet structures, where $\alpha\in\{x,y\}$ indicates whether the path $\mathcal{C}$ runs along the $k_{x}$ or $k_{y}$ direction. Specifically, a non-vanishing invariant $m_{\alpha}$ corresponds to a non-trivial permutation $\pi_{\alpha}$ of the two energies, that is, exchanging the eigenvalues $\epsilon_{1}$ and $\epsilon_{2}$ of the two-sheeted Riemann surface:
$m_{\alpha}=1-\delta_{\pi_{\alpha}(1),2}\,.$ (4)
This establishes $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ topological order in the non-Hermitian spectrum. Based on the four states distinguished by the combinations of topologically protected closed Fermi cuts, we construct the analogy to the ground-state manifold of the toric code in an exemplary model. Any closed Fermi cut is mapped to a non-contractible loop of flipped spins in the toric code, and the four different ground states are represented by $(m_{x},m_{y})\in\{(0,0),(1,0),(0,1),(1,1)\}$. The ground-state manifold is realized by the tunable non-Hermitian system with the Bloch Hamiltonian
@@
At the same time, added bands affect the ground-state manifold of the non-Hermitian multi-band model, which comprises any fully non-degenerate spectral structure. The number of topologically distinguishable Riemann surface structures increases, because additional non-contractible Fermi cuts along the $k_{x}$ and $k_{y}$ direction are possible between any two bands. However, intersecting Fermi cuts induce EP3 excitations if one energy sheet is part of both cuts. Therefore not all combinations of Fermi cuts are part of the ground-state manifold. A set of topological invariants can be defined for each ground state by generalizing Eq. (4) to all pairs $(p,q)$ of energy sheets and repeated application of the permutation $\pi_{\alpha}$:
$m_{\alpha}^{pq}=1-\delta\Bigl{[}\,\sum_{j=1}^{n-1}\delta\left[p-\pi_{\alpha}^{j}(q)\right]\Bigr{]}\,,$ (7)
where $\delta[n]$ denotes the Kronecker delta $\delta_{0,n}$. This establishes the ground-state manifold as a subspace of the space with $\left(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\right)^{m}$ topological order, where $m=\sum_{j=1}^{n-1}j=\tfrac{1}{2}(n-1)n$.
@@
Implementation on metasurfaces and in single-photon interferometry.—The implementation of different topologically protected states in the non-Hermitian toric-code analog requires a high degree of parametric tunability over a toroidal parameter space. Measuring the connectedness of the Riemann surface structure on the other hand is achievable, for example by sampling the real part of the spectrum along non-contractible loops in the Brillouin zone. We remark, that conclusions about the connectedness based state evolution are not possible, since the adiabatic theorem does not hold for the non-Hermitian models discussed here [35, 36, 37]. Prominent platforms providing the necessary tunability are optical or plasmonic metasurfaces, which rely on artificial units cells defined in periodic arrays of nanoantennas [38, 39]. Loss generates non-Hermitian contributions to the dynamics of these systems. By carefully adjusting the shape, spacing and orientation of the antennas in the surrounding medium or on the host surface, the parameters of the system can be precisely controlled. Such setups are capable of realizing the two-band model $H_{\text{FC}}(\bm{k})$ in a two-orbital square lattice given full control over the onsite terms and hoppings up to next-nearest-neighbor distance. Mechanical metamaterials, in which individual oscillators are driven depending on the state of the system, provide another feasible platform. Here, the dynamical control of the oscillator response allows for the observation of EP pair creation and their threading through the Brillouin zone. Thus these materials are capable of resolving the transition between topologically distinct states. Single-photon interferometry was recently established as a different platform for the study of non-Hermitian topological features [40]. These newly designed interferometers can, for example, encode the evolution of a non-Hermitian three-band Bloch Hamiltonian and realize a braid protected EP3 [40, 30]. By instead encoding the parameters of $H_{\text{FC}}$, the non-Hermitian analog of the toric-code ground states can be realized in this state-of-the-art platform.
@@
Acknowledgments.—We are grateful to Joost Slingerland for his lectures during the Young Researchers School 2024 in Maynooth and to Lukas König for fruitful discussion. We acknowledge funding from the Max Planck Society’s Lise Meitner Excellence Program 2.0.